OpenStudy (anonymous):

Find equation of a plane which is equidistant to parallel planes 2x-2y-z+7=0 and 8x-8y-4z+1=0.

OpenStudy (anonymous):

express 2º equation like this: 2x-2y-1z=-1/4 and 2º 2x-2y-z=-7 so, notice that bouth planes are at the "same side" of the origin (bouth have negative distance) so just rest 1/4 from 7, to get ?

OpenStudy (anonymous):

got it @Valdas ?

OpenStudy (anonymous):

What do you mean by "rest 1/4 from 7, to get ?"?

OpenStudy (anonymous):

\[7-\frac{1}{4}=\frac{27}{4}\]

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

And what does that number represent? The distance between those planes?

OpenStudy (anonymous):

it is the distance that you need in the plane formula.

OpenStudy (anonymous):

2x-2y-1z=-27/4

OpenStudy (amistre64):

since the planes are parallel, i would find a point in one plane, and anchor the normal to it to form a line; i would then determine where that line punctures the other plane at. this gives two points to deduce a midpoint with; that midpoint would be used to define the plane between them

OpenStudy (anonymous):

to complicated

OpenStudy (amistre64):

im a complicated person :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

little mistake: 27/4 is the distance between planes, what you need is take half of it and add it to 1/4

OpenStudy (anonymous):

so 27/8+1/4=29/8

OpenStudy (anonymous):

and plane formula: 2x-2y-1z=-29/8

OpenStudy (anonymous):

Distance between original planes is 9/4 (right?), so the distance between any those planes and the new one should be 9/8. But The distance between 2x-2y-z+7=0 and 2x-2y-1z=-29/8 is 3/8 if I did all my calculations correctly

OpenStudy (anonymous):

distance between original planes is 27/4

OpenStudy (anonymous):

This is how I calculated the distance. I took point (0,0,7) from plane 2x-2y-z+7=0 and found distance between this point and 8x-8y-4z+1=0 using distance formula. \[\frac{ |8*0-8*0-4*7+1| }{ \sqrt{64+64+16} }=\frac{ 27 }{ 12 }=\frac{ 9 }{ 4 }\]

OpenStudy (amistre64):



OpenStudy (amistre64):

as myko points out; we can determine the needed point by taking the average of the points where they cross the z axis

OpenStudy (amistre64):

x,y = 0 2x-2y-z+7=0 0 0 -------------- -z+7=0 ; z=7 8x-8y-4z+1=0 0 0 --------------- -4z+1=0 ; z=1/4 (7+1/4)/2 = 29/8

OpenStudy (amistre64):

use the point (0,0,29/8) and the normal <2,-2,-1> as your midplane

OpenStudy (anonymous):

Is the way I calculated the distance is somehow wrong?

OpenStudy (amistre64):

the more complicated we try to do things, the more room for error creeps in :) if you found the perp distance, that would be fine as well

OpenStudy (amistre64):

the mathing looks fine; so, 9/4 has to be applied to the line setup

OpenStudy (amistre64):

or rather 9/4 divided by 2 is 9/8 for the half distance

OpenStudy (anonymous):

So, as I understand, the distance between 2x-2y-z+7=0 and 2x-2y-1z+29/8=0 should be equal to 9/8? But I get, that it's 3/8 \[\frac{ |2*0-2*0-7+29/8| }{ \sqrt{4+4+1} }=\frac{ 27/8 }{ 9 }=\frac{ 3 }{ 8 }\]

OpenStudy (amistre64):

2x-2y-z+7=0 the line is: x = 0 + 2t y = 0 - 2t z = 7 - t using the other plane and these defs 8x-8y-4z +1=0 2t -2t (7-t) --------------- -28+4t+1=0 -27+4t =0 ; t = 27/4 is the distance between the planes in order for the line to pierce the other plane; half the distance defines it midpoint. t=27/8 x = 0 + 2t; x=27/4 y = 0 - 2t; y=-27/4 z = 7 - t; z= 29/8

OpenStudy (anonymous):

I did a mistake in my calculation. The distance is correct. Thanks for help

OpenStudy (amistre64):

;) good luck